OpenStudy (anonymous):

step by step factor x^3 - x^2 - 4x + 4

OpenStudy (anonymous):

Use the order of operations

OpenStudy (luigi0210):

you could also use grouping

OpenStudy (anonymous):

ok not sure how that works with so many terms is why i am asking how to factor this

OpenStudy (anonymous):

Or you could find one factor in that equation then use synthetic division to solve it

OpenStudy (anonymous):

looking for real zeroes and yand x ints and end behavior

OpenStudy (anonymous):

ty for assistance just dont have my brain wrapped around this

OpenStudy (tkhunny):

1) It's a cubic expression. It will only be luck if you can factor it. 2) Look around for common factors in various chunks. x^2 is common in the first two terms. -4 is common in the last two terms. Factor those and reassess.

OpenStudy (anonymous):

I guess I would group the first two and the last two terms?

OpenStudy (tkhunny):

That does look promising. Let's see what you get.

OpenStudy (anonymous):

ok going to try

OpenStudy (anonymous):

Find the factors of the constant and the factors of the leading coeffiecient. So you'll have +/- 1, 2, 4 so plug one of those factors in the place of the x and see what factors out to =0 and once you have that you can use synthetic division to find out your other factors. Even though they are already given to you once you have the factors of 4 since there are only 3 factors and the leading exponent is saying there are 3 possible factors. Hope this helps, idk what level of math you're at though.

OpenStudy (anonymous):

college algebra just started learning synthetic division

OpenStudy (anonymous):

You didn't learn this in high school?

OpenStudy (tkhunny):

No need. The problem statement WANTS you to factor it. Just stick with the grouping.

OpenStudy (anonymous):

when I graduated in 1983 ( Not to date myself LOL ) it wasn't required.

OpenStudy (anonymous):

Oh haha got ya, yeah not something you really need to remember in the real world but it's super easy anyways yeah just factor it simply and you got it, it's not a hard problem c:

OpenStudy (anonymous):

is for me

OpenStudy (anonymous):

I got x^2 ( X-1) - 4 (x+1)

OpenStudy (anonymous):

I know I need to go further but how

OpenStudy (tkhunny):

Close. The last one should be (x-1). Try that part again.

OpenStudy (anonymous):

just saw that : )

OpenStudy (tkhunny):

More factoring. You;ve a common factor of (x-1) in both of those remaining terms.

OpenStudy (anonymous):

so that (x-1) ^2

OpenStudy (anonymous):

would that be x^2 - 4 (x-1)^2 so far

OpenStudy (tkhunny):

How did you get that? \(x^{2}(x-1) - 4(x-1) = (x^{2} - 4)(x-1)\)

OpenStudy (tkhunny):

The \(x^{2}\) didn't magically square when you factored it out of the first two terms. The \(-4\) didn't magically square when you factored it out of the first two terms. Why would the \((x-1)\) square when you factored it out of the two remaining terms?

OpenStudy (anonymous):

I shouldnt have combined the terms lol

OpenStudy (tkhunny):

You have to get this property stuck in your brain. \(a\cdot(b+c) = ab + ac = ba + ca = (b+c)\cdot a\) It works both ways.

OpenStudy (anonymous):

I set it up like this (x^3 - x^2) - (4x+4)

OpenStudy (anonymous):

then pulled out x^2 from the first set

OpenStudy (anonymous):

then 4 from the second set

OpenStudy (anonymous):

leaving me with x^2 (x-1) - 4(x-1)

OpenStudy (anonymous):

Okai you're making it harder than it is. first things first. The highest exponent is telling you how many factors there are going to be. So there are 3 factors. you take the factors of the constant which is 4. Then you take the factors of the leading coefficient. In this case that is 1. so your factors are +/- 1,2,4. Try one of the factors. I chose 1. so the problem will look like this (1)^3-(1)^2-4(1)+4 that equals 1-1-4+4=0 so there you have your first factor.

OpenStudy (anonymous):

and I am lost from tht point

OpenStudy (anonymous):

Then you're going to use synthetic division which is gonna look like this...